# Supervised heterogeneous network estimation via survival-based Bayesian graphical models

### v1

@@ 3 Simulation
Simulation studies are conducted to evaluate the performance of the proposed approach under the following settings. (a) $K=2$ and $3$ and $p=100$. (b) Consider two settings for the sample size. One is a balanced design where there are 150 subjects in each subgroup, while the other is an imbalanced design with sample sizes being $100$ and $200$ for $K=2$ and $100,150$, and $200$ for $K=3$. Thus, there are $n=300$ subjects for $K=2$, and $n=450$ for $K=3$ in total. (c) The predictor observations $\left\{\bm{x}_{i}\right\}$ in the $k$th subgroup are simulated from $\mathcal{N}\left(\bm{0},\bm{\Omega}_{k}^{-1}\right)$, where $\bm{\Omega}_{k}$ is generated based on the corresponding network structure. Specifically, following [6], we simulate each network with ten unconnected subnetworks. To comprehensively examine the performance of the proposed approach under different degrees of similarity across subgroup-specific precision matrices, we consider three specific settings S1-S3, with the number of subnetworks sharing the same sparsity structure in all subgroups being 3, 5, and 7, respectively. For each subnetwork, we consider three types of structure, including the Power-law network with the degrees following a power-law distribution, the Nearest-neighbor network with $p/10$ points being first generated on a unit square and one nearest neighbor then being identified for each point by calculating the pairwise distances of the $p/10\times(p/10-1)/2$ points, and the Erdös-Rényi network with edges randomly added with a probability of 0.1. For the $k$th subgroup, set ${\omega}_{k,jj}=0$ and ${\omega}_{k,jl}=0$ if $j$ and $l(\neq j)$ are not connected in the $k$th simulated network; otherwise, the signs of ${\omega}_{k,jl}$’s are randomly set to positive or negative with an equal probability, and their absolute values are independently generated from uniform distributions with support on $[0.5,1]$ for $K=2$, and $[0.8,1]\times I\left\{k=1\right\}+[1.8,2]\times I\left\{k=2\right\}+[2.8,3]\times I\left\{k=3\right\}$ for $K=3$. We further add $0.01+\left|\lambda_{\min}\left(\bm{\Omega}_{k}\right)\right|$ to the diagonal elements of the $k$th precision matrix to ensure that $\bm{\Omega}_{k}$ is positive definite. (d) Each logarithmic survival observation $t_{i}$ is generated as follows. First, the logarithmic survival times of the $k$th subgroup are generated from $\mathcal{N}\left(\bm{\beta}_{k}^{\top}\bm{x}_{i},0.01^{2}\right)$, where five elements of $\bm{\beta}_{k}$ are non-zero, and the rest $p-5$ elements are all zero with $\bm{\beta}_{1}=\left(2,2,2,2,2,0,\ldots,0\right)^{\top}$, $\bm{\beta}_{2}=-\left(2,2,2,2,2,0,\ldots,0\right)^{\top}$, and $\bm{\beta}_{3}=\left(0,0,1,1,1,1,1,0,\ldots,0\right)^{\top}$, respectively. Then, the censoring time is simulated from a Gamma distribution with the shape and scale parameters controlling the censoring rate to be $\sim 20\%$. There are a total of 36 scenarios.
@@
In addition to the proposed approach, six alternative approaches are also considered. BCGS is a Bayesian approach for simultaneously clustering and learning multiple sparse graphical models using a joint spike-and-slab graphical lasso prior [5]. SCAN is a mixture Gaussian graphical model-based approach that applies the group lasso penalty to accommodate common structures across unknown heterogeneous subgroups [11]. TGemBag and TJGL are two “ideal approaches” which assume that the true subgroup memberships are known and adopt GemBag and JGL for jointly estimating multiple networks for the true subgroups, respectively. Here, GemBag is a Bayesian graphical model for accommodating similarity as well as difference of multiple networks [26], and JGL employs a likelihood-based joint graphical Lasso for multiple network estimations [6]. FMGemBag and FMJGL are two two-stage approaches, where in the first stage, a finite mixture (FM) of AFT models developed in [22] is used to obtain the subgroup memberships of all subjects, which can account for right censoring and intangible heterogeneous predictor-survival associations, and in the second stage, GemBag and JGL are adopted for network estimation based on the estimated subgroup memberships, respectively. Among these methods, BCGS, SCAN, TJGL, and FMJGL can be realized using R package JGL, and TGemBag and FMGemBag can be realized using R codes at https://github.com/xinming104/GemBag. BCGS and SCAN achieve identification of unknown heterogeneity and network structures simultaneously, however, are unsupervised and only account for predictors. GemBag and JGL are developed for estimating precision matrices with known subgroup memberships. We consider two types of subgroup memberships for GemBag and JGL, including the one based on the FM approach, which utilizes the information of survival for heterogeneous analysis, and the “ideal” type based on the true subgroup memberships.
@@
To evaluate the performance of different approaches, the following measures are calculated: (1) Clustering error (CE) for evaluating the heterogeneity identification performance, defined as $\operatorname{CE}(\hat{\varphi},\varphi^{*})\equiv\left(\begin{array}[]{l}n\\ 2\end{array}\right)^{-1}\left|\left\{(i,j):{I}\left\{\hat{\varphi}_{i}=\hat{\varphi}_{j}\right\}\neq{I}\left\{\varphi_{i}^{*}=\varphi_{j}^{*}\right\};i<j\right\}\right|$ with $\hat{\varphi}$ and ${\varphi}^{*}$ being the estimated and true subgroup memberships. (2) Precision matrix squared error (PME) for evaluating the estimation performance, defined as $\operatorname{PME}=\sum_{k=1}^{K}\frac{\left\|\widetilde{\bm{\Omega}}_{k}-\bm{\Omega}_{k}^{*}\right\|_{F}}{K}.$ (3) True and false positive rates (TPR and FPR) for evaluating the network identification performance, defined as $\operatorname{TPR}=\sum_{k=1}^{K}\frac{\sum_{j<l}{I}\left\{\omega_{k,jl}^{*}\neq 0,\widetilde{\omega}_{k,jl}\neq 0\right\}}{K\sum_{j<l}{I}{\left\{\omega_{k,jl}^{*}\neq 0\right\}}}$ and $\operatorname{FPR}=\sum_{k=1}^{K}\frac{\sum_{j<l}{I}{\left\{{\omega}_{k,jl}^{*}=0,\widetilde{\omega}_{k,jl}\neq 0\right\}}}{K\sum_{j<l}{I}\left\{{\omega}_{k,jl}^{*}=0\right\}}.$ Here, similar to other network-based heterogeneity analysis studies [11, 5], we focus on the performance of estimation and identification on $\omega_{k,jl}$’s, which can be more challenging than on $\mu_{k,j}$’s and $\beta_{k,j}$’s.
@@
100 replicates are conducted for each scenario, where the true value of $K$ is used for all approaches. The summary results for $K=2$ are shown in Tables 1-3, and the rest of the results are provided in Supplementary Materials. Under all simulated scenarios, the proposed approach consistently demonstrates superior or competitive performance compared to the alternative approaches. In terms of subgroup identification, the proposed approach performs significantly better with much smaller CEs. Among the alternatives, the FM methods tend to exhibit the poorest subgroup identification performance, while SCAN and BCGS have better subgroup identification accuracy but are still much inferior compared to the proposed approach.
@@
CE 0.012(0.010) 0.158(0.174) 0.162(0.176) - - 0.274(0.120) 0.274(0.120)
PME 4.450(0.374) 7.160(4.339) 7.243(4.615) 4.958(1.616) 8.249(1.091) 6.989(1.190) 7.981(0.816)
TPR 0.933(0.031) 0.775(0.192) 0.765(0.194) 0.817(0.171) 0.886(0.099) 0.877(0.089) 0.920(0.044)
FPR 0.023(0.007) 0.033(0.010) 0.032(0.009) 0.021(0.009) 0.025(0.012) 0.115(0.020) 0.083(0.017)
Additionally, to gain a deeper understanding of the similarity-based prior, we illustrate a representative replicate under the scenario with setting S3 and an imbalanced design. The heatmaps depicting the true sparsity structures and the estimated results of the precision matrices with our proposed approach, SCAN, and TGemBag, are presented in Supplementary Figures S1 and S2. The findings indicate that our proposed approach excels in accurately identifying true sparsity structures and more effectively accommodating the similarity structures across different networks.
@@
CE 0.027(0.031) 0.142(0.149) 0.145(0.150) - - 0.265(0.116) 0.265(0.116)
PME 3.417(0.241) 5.082(2.511) 5.170(2.652) 3.072(0.373) 5.007(0.624) 4.896(0.902) 5.317(0.813)
TPR 0.963(0.027) 0.825(0.163) 0.812(0.162) 0.930(0.058) 0.958(0.036) 0.940(0.054) 0.947(0.060)
FPR 0.024(0.008) 0.031(0.009) 0.030(0.009) 0.027(0.008) 0.026(0.010) 0.131(0.033) 0.076(0.028)
@@ 4 Data Analysis
We illustrate the proposed approach in constructing gene networks of different subgroups of cutaneous melanoma (SKCM) from The Cancer Genome Atlas (TCGA) data. Date is downloaded from the TCGA Provisional using R package cgdsr. We consider the metastatic subjects and use the mRNA gene expression data and overall survival time for the predictors and response, respectively. $360$ subjects are available with $19,039$ gene expression measurements. Among them, $53.6\%$ are non-censored, with the observed survival times ranging from 2.6 to 357.1 months (median 44.05 months). As the sample size is relatively small, we follow the existing studies [1] and conduct gene filtering. Specifically, we first identify 2,260 genes with significance levels below $0.01$ from univariate Cox-regression models and then focus on those present in six pathways (pyrimidine metabolism, P53 signaling, cysteine and methionine metabolism, chemokine signaling, cell cycle, and hedgehog signaling pathways) that have demonstrated important implications for SKCM in the literature. With this preprocessing, 91 genes are obtained for downstream analysis.
@@
As the number of subgroups $K$ is unknown, we consider the candidate set $\left\{2,3,4,5,6,7\right\}$ and adopt the Gap statistic for selecting the optimal one. With the proposed approach, the Gap statistic leads to two subgroups with sizes of 128 (subgroup 1) and 232 (subgroup 2). The two networks consist of 181 and 345 edges among 80 and 83 genes, respectively. Among them, 52 edges are common and all have the same signs. A graphical representation is provided in Figure 3, and more detailed estimation results are provided in Supplementary Materials. To gain deeper insights into the identified networks, we conduct Gene Ontology (GO) enrichment analysis using DAVID 2021 to examine the functional and biological connections of the related genes. Some significantly enriched GO terms are obtained for the common edges, and genes that exhibit unique edges in the two subgroups are found to be linked to different GO terms. The sensible biological findings provide support to the validity of the proposed network estimation. We refer to Supplementary Materials for the detailed GO enrichment analysis results.
Table 3: Simulation results for the nearest-neighbor network under the scenarios with $K=2$. In each cell, mean (SD) based on 100 replicates.
proposed BCGS SCAN TGemBag TJGL FMGemBag FMJGL
S1 with the balanced design
CE 0.001(0.003) 0.001(0.003) 0.001(0.003) - - 0.214(0.047) 0.214(0.047)
PME 3.867(0.225) 4.051(0.193) 3.996(0.194) 2.827(0.254) 5.482(0.352) 7.902(2.318) 8.074(1.631)
TPR 0.990(0.010) 0.989(0.013) 0.988(0.014) 0.989(0.012) 0.988(0.013) 0.818(0.135) 0.817(0.137)
FPR 0.050(0.013) 0.059(0.017) 0.057(0.016) 0.050(0.014) 0.051(0.016) 0.222(0.031) 0.149(0.041)
S2 with the balanced design
CE 0.006(0.010) 0.007(0.009) 0.007(0.009) - - 0.212(0.065) 0.212(0.065)
PME 3.867(0.204) 4.040(0.150) 3.986(0.144) 2.810(0.219) 5.471(0.346) 5.705(1.986) 6.338(1.437)
TPR 0.993(0.007) 0.991(0.011) 0.989(0.012) 0.992(0.009) 0.990(0.010) 0.937(0.089) 0.932(0.094)
FPR 0.049(0.012) 0.057(0.016) 0.055(0.016) 0.049(0.014) 0.049(0.015) 0.209(0.025) 0.135(0.032)
S3 with the balanced design
CE 0.024(0.025) 0.091(0.169) 0.081(0.159) - - 0.225(0.094) 0.225(0.094)
PME 3.618(0.186) 5.639(4.789) 4.951(2.931) 2.816(0.211) 5.504(0.381) 5.865(1.893) 7.353(5.279)
TPR 0.989(0.011) 0.923(0.169) 0.931(0.158) 0.992(0.008) 0.988(0.011) 0.910(0.114) 0.926(0.070)
FPR 0.046(0.013) 0.057(0.016) 0.055(0.015) 0.048(0.014) 0.048(0.016) 0.201(0.027) 0.129(0.033)
S1 with the imbalanced design
CE 0.002(0.004) 0.003(0.006) 0.003(0.006) - - 0.222(0.042) 0.222(0.042)
PME 4.051(0.259) 4.247(0.194) 4.211(0.207) 3.078(0.298) 5.507(0.347) 8.231(1.964) 8.267(1.386)
TPR 0.971(0.024) 0.961(0.030) 0.957(0.033) 0.964(0.027) 0.986(0.017) 0.796(0.117) 0.802(0.114)
FPR 0.050(0.013) 0.058(0.016) 0.056(0.016) 0.050(0.014) 0.057(0.016) 0.218(0.036) 0.154(0.041)
S2 with the imbalanced design
CE 0.006(0.014) 0.009(0.015) 0.009(0.015) - - 0.228(0.066) 0.228(0.066)
PME 4.032(0.234) 4.253(0.208) 4.216(0.212) 3.057(0.265) 5.490(0.344) 6.574(1.815) 6.957(1.285)
TPR 0.979(0.013) 0.963(0.026) 0.959(0.028) 0.968(0.020) 0.987(0.013) 0.899(0.090) 0.897(0.092)
FPR 0.049(0.013) 0.056(0.017) 0.055(0.016) 0.049(0.015) 0.055(0.016) 0.213(0.032) 0.149(0.037)
S3 with the imbalanced design
CE 0.022(0.024) 0.067(0.108) 0.067(0.109) - - 0.238(0.051) 0.238(0.051)
PME 3.750(0.208) 4.688(1.432) 4.699(1.554) 3.026(0.282) 5.527(0.395) 5.656(1.355) 6.219(1.098)
TPR 0.974(0.018) 0.923(0.121) 0.917(0.121) 0.966(0.029) 0.986(0.012) 0.923(0.062) 0.934(0.058)
FPR 0.045(0.014) 0.056(0.016) 0.054(0.016) 0.048(0.015) 0.054(0.016) 0.199(0.027) 0.142(0.034)
@@
Comparative analysis using the alternative approaches is also carried out with $K=2$. The summary results are reported in Table S4 of Supplementary Materials, which provide the number of subjects and network edges in the two subgroups identified by different approaches and their overlaps. The subgroups identified using different approaches are matched by correlation. In Table S4, different approaches are observed to identify subgroups with different subjects. To get a deeper understanding of the heterogeneity analysis performance, we examine whether these subgroups are associated with survival time. In Figure 4, we present the Kaplan–Meier (KM) curves of survival time of different subgroups identified by the proposed and alternative approaches. It is observed that the KM curves of the two subgroups are significantly distinguished from each other for the proposed approach (P value: 0.005). Comparatively, the FM approach shows less significant differences. BCGS and SCAN approaches fail to identify subgroups with distinct survival. These biologically reasonable findings indicate that the proposed approach can satisfactorily identify survival-related subgroups and also provide support for its effectiveness in clinical subgroup identification.
In Table S4, different approaches demonstrate a moderate number of overlapping edges. To provide indirect support for the network estimation results, we use the resampling strategy and randomly divide the data into a training and a testing set. As in the literature [7], the values of the negative log-likelihood statistic (NLS) for the 91 genes are calculated to evaluate prediction accuracy. The average NLS values of 100 resamplings are 7102.118 for the proposed approach, 7658.492 for BCGS, 7674.739 for SCAN, 8044.482 for FMGemBag, and 8221.576 for FMJGL, which suggest a satisfactory network estimation accuracy of the proposed approach.

### v2

@@ 3 Simulation
Simulation studies are conducted to evaluate the performance of the proposed approach under the following settings. (a) $K=2$ and $3$ and $p=100$. (b) Consider two settings for the sample size. One is a balanced design where there are 150 subjects in each subgroup, while the other is an imbalanced design with sample sizes being $100$ and $200$ for $K=2$ and $100,150$, and $200$ for $K=3$. Thus, there are $n=300$ subjects for $K=2$, and $n=450$ for $K=3$ in total. (c) The predictor observations $\left\{\bm{x}_{i}\right\}$ in the $k$th subgroup are simulated from $\mathcal{N}\left(\bm{0},\bm{\Omega}_{k}^{-1}\right)$, where $\bm{\Omega}_{k}$ is generated based on the corresponding network structure. Specifically, following [7], we simulate each network with ten unconnected subnetworks. To comprehensively examine the performance of the proposed approach under different degrees of similarity across subgroup-specific precision matrices, we consider three specific settings S1-S3, with the number of subnetworks sharing the same sparsity structure in all subgroups being 3, 5, and 7, respectively. For each subnetwork, we consider three types of structure, including the power-law, nearest-neighbor, and Erdös-Rényi networks. The detailed settings are provided in Section S6 of the supplementary materials. (d) Each logarithmic survival observation $t_{i}$ is generated as follows. First, the logarithmic survival times of the $k$th subgroup are generated from $\mathcal{N}\left(\bm{\beta}_{k}^{\top}\bm{x}_{i},0.01^{2}\right)$, where five elements of $\bm{\beta}_{k}$ are non-zero, and the rest $p-5$ elements are all zero with $\bm{\beta}_{1}=\left(2,2,2,2,2,0,\ldots,0\right)^{\top}$, $\bm{\beta}_{2}=-\left(2,2,2,2,2,0,\ldots,0\right)^{\top}$, and $\bm{\beta}_{3}=\left(0,0,1,1,1,1,1,0,\ldots,0\right)^{\top}$, respectively. Then, the censoring time is simulated from a Gamma distribution with the shape and scale parameters controlling the censoring rate to be $\sim 20\%$. There are a total of 36 scenarios.
@@
In addition to the proposed approach, six alternative approaches are also considered. BCGS is a Bayesian approach for simultaneously clustering and learning multiple sparse graphical models using a joint spike-and-slab graphical lasso prior [6]. SCAN is a mixture Gaussian graphical model-based approach that applies the group lasso penalty to accommodate common structures across unknown heterogeneous subgroups [12]. TGemBag and TJGL are two “ideal approaches” which assume that the true subgroup memberships are known and adopt GemBag and JGL for jointly estimating multiple networks for the true subgroups, respectively. Here, GemBag is a Bayesian graphical model for accommodating similarity as well as difference of multiple networks [28], and JGL employs a likelihood-based joint graphical Lasso for multiple network estimations [7]. FMGemBag and FMJGL are two two-stage approaches, where in the first stage, a finite mixture (FM) of AFT models developed in [24] is used to obtain the subgroup memberships of all subjects, which can account for right censoring and unknown heterogeneous predictor-survival associations, and in the second stage, GemBag and JGL are adopted for network estimation based on the estimated subgroup memberships, respectively. Among these methods, BCGS, SCAN, TJGL, and FMJGL can be realized using or extending the functionalities provided by the R package JGL, and TGemBag and FMGemBag can be realized using R codes at https://github.com/xinming104/GemBag. BCGS and SCAN achieve identification of unknown heterogeneity and network structures simultaneously, however, are unsupervised and only account for predictors. GemBag and JGL are developed for estimating precision matrices with known subgroup memberships. We consider two types of subgroup memberships for GemBag and JGL, including the one based on the FM approach, which utilizes the information of survival for heterogeneous analysis, and the “ideal” type based on the true subgroup memberships.
@@
To evaluate the performance of different approaches, we consider the clustering error (CE) for evaluating the heterogeneity identification performance, precision matrix squared error (PME) for evaluating the estimation performance, and true and false positive rates (TPR and FPR) for evaluating the network identification performance. The details are provided in Section S6 of the supplementary materials. Here, similar to other network-based heterogeneity analysis studies [12], we focus on the performance of estimation and identification on $\omega_{k,jl}$’s, which can be more challenging than on $\mu_{k,j}$’s and $\beta_{k,j}$’s.
@@
100 replicates are conducted for each scenario, where the true value of $K$ is used for all approaches. The summary results for $K=2$ under power-law and Erdös-Rényi networks are presented in Tables 2-3, and the rest of the results are provided in Section S6 of the supplementary materials. Under all simulated scenarios, the proposed approach consistently demonstrates superior or competitive performance compared to the alternative approaches. In terms of subgroup identification, the proposed approach performs significantly better with much smaller CEs. Among the alternatives, the FM methods tend to exhibit the poorest subgroup identification performance, while SCAN and BCGS have better subgroup identification accuracy but are still much inferior compared to the proposed approach.
@@
CE 0.013(0.009) 0.158(0.174) 0.162(0.176) - - 0.274(0.120) 0.274(0.120)
PME 4.449(0.381) 7.160(4.339) 7.243(4.615) 4.958(1.616) 8.249(1.091) 6.989(1.190) 7.981(0.816)
TPR 0.932(0.030) 0.775(0.192) 0.765(0.194) 0.817(0.171) 0.886(0.099) 0.877(0.089) 0.920(0.044)
FPR 0.022(0.006) 0.033(0.010) 0.032(0.009) 0.021(0.009) 0.025(0.012) 0.115(0.020) 0.083(0.017)
@@
CE 0.028(0.030) 0.142(0.149) 0.145(0.150) - - 0.265(0.116) 0.265(0.116)
PME 3.416(0.241) 5.082(2.511) 5.170(2.652) 3.072(0.373) 5.007(0.624) 4.896(0.902) 5.317(0.813)
TPR 0.962(0.026) 0.825(0.163) 0.812(0.162) 0.930(0.058) 0.958(0.036) 0.940(0.054) 0.947(0.060)
FPR 0.023(0.007) 0.031(0.009) 0.030(0.009) 0.027(0.008) 0.026(0.010) 0.131(0.033) 0.076(0.028)
Additionally, to gain a deeper understanding of the similarity-based prior, we illustrate a representative replicate under the scenario with setting S3 and an imbalanced design. The heatmaps depicting the true sparsity structures and the estimated results of the precision matrices with our proposed approach, SCAN, and TGemBag, are presented in Figures S3-S4 (supplementary materials). The findings indicate that our proposed approach excels in accurately identifying true sparsity structures and more effectively accommodating the similarity structures across different networks.
Following previous studies [11, 12, 31], the main simulations assume that the number of subgroups $K$ is known for a more fair comparison. When $K$ is unknown, we recommend using the BIC for model selection. As an illustration, we consider scenarios with a power-law network and balanced design and evaluate the empirical performance of BIC over 100 replicates with candidate values $K=1,\ldots,5$. The selection frequencies, summarized in Table S5 (supplementary materials), indicate that BIC generally performs well. In particular, the S3 configuration, which exhibits greater subgroup similarity than S1 and S2, poses more challenges for accurate identification. For the BIC-selected values of $K$, Table S6 (supplementary material) reports the corresponding identification and estimation results. In cases where $K$ is incorrectly estimated, label alignment is performed minimizing the estimation error of $\bm{\Psi}$ [22]. The results remain consistent with those presented in Table 2 and Table S2 (supplementary materials).
@@ 4 Data Analysis
We illustrate the proposed approach in constructing gene networks of different subgroups of cutaneous melanoma (SKCM) from The Cancer Genome Atlas (TCGA) data. Following existing literature [1], we consider the metastatic subjects and use the mRNA gene expression data and overall survival time for the predictors and response, respectively. $360$ subjects are available with $19,039$ gene expression measurements. Among them, $53.6\%$ are non-censored, with the observed survival times ranging from 2.6 to 357.1 months (median 44.05 months). As the sample size is relatively small, we follow the existing studies [12] and conduct gene filtering. Specifically, we first identify 2,260 genes with significance levels below $0.01$ from univariate Cox-regression models and then focus on those present in six pathways (pyrimidine metabolism, P53 signaling, cysteine and methionine metabolism, chemokine signaling, cell cycle, and hedgehog signaling pathways) that have demonstrated important implications for SKCM in the literature. With this preprocessing, 91 genes are obtained for downstream analysis.
@@
As the number of subgroups $K$ is unknown, we consider the candidate set $\left\{1,2,3,4,5,6,7\right\}$ and adopt the BIC criterion for selecting the optimal one. With the proposed approach, the BIC criterion leads to two subgroups with sizes of 128 (subgroup 1) and 232 (subgroup 2). The two networks consist of 181 and 345 edges among 80 and 83 genes, respectively. Among them, 52 edges are common and all have the same signs. A graphical representation is provided in Figure 1, and more detailed estimation results are provided in the supplementary materials. To gain deeper insights into the identified networks, we conduct Gene Ontology (GO) enrichment analysis using DAVID 2021 to examine the functional and biological connections of the related genes. Some significantly enriched GO terms are obtained for the common edges, and genes that exhibit unique edges in the two subgroups are found to be linked to different GO terms. The sensible biological findings provide support to the validity of the proposed network estimation. We refer to Section S7 of the supplementary materials for the detailed GO enrichment analysis results.
@@
Comparative analysis using the alternative approaches is also carried out with $K=2$. The summary results are reported in Table S7 (supplementary materials), which provide the number of subjects and network edges in the two subgroups identified by different approaches and their overlaps. The subgroups identified using different approaches are matched by correlation. In Table S7 (supplementary materials), different approaches are observed to identify subgroups with different subjects. To get a deeper understanding of the heterogeneity analysis performance, we examine whether these subgroups are associated with survival time. In Figure 2, we present the Kaplan–Meier (KM) curves of survival time of different subgroups identified by the proposed and alternative approaches. It is observed that the KM curves of the two subgroups are significantly distinguished from each other for the proposed approach (P value: 0.005). Comparatively, the FM approach shows less significant differences. BCGS and SCAN approaches fail to identify subgroups with distinct survival. These biologically reasonable findings indicate that the proposed approach can satisfactorily identify survival-related subgroups and also provide support for its effectiveness in clinical subgroup identification.
@@
In Table S7 (supplementary materials), different approaches demonstrate a moderate number of overlapping edges. To provide indirect support for the network estimation results, we use the resampling strategy and randomly divide the data into a training and a testing set. As in the literature [9], the values of the negative log-likelihood statistic (NLS) for the 91 genes are calculated to evaluate prediction accuracy. The average NLS values of 100 resamplings are 7102.118 for the proposed approach, 7658.492 for BCGS, 7674.739 for SCAN, 8044.482 for FMGemBag, and 8221.576 for FMJGL, which suggest a satisfactory network estimation accuracy of the proposed approach.